# They See Me Rolling: High-Speed Event Vision-Based Tactile Roller Sensor for Large Surface Inspection

### v2

@@ V-A Experimental Setup
The subsequent Results section (Sec. VI) details the outcomes of these experiments. The evaluations focus on: (1) 3D reconstruction accuracy, primarily quantified using Mean Absolute Error (MAE) against ground truth geometries, including an analysis of the proposed BMA fusion strategy; (2) the relationship between rolling speed and reconstruction accuracy; (3) the effectiveness of large surface reconstruction through measurement stitching; and (4) the speed and accuracy achieved during high-speed Braille character recognition. The MAE is calculated by comparing the reconstructed depth map, $Z_{f}$, directly to the ground truth depth map, $Z_{gt}$. The ground truth map is created by first identifying the contact patch in the reconstructed data and then using the known physical geometry of the object to compute the ideal depth value for each pixel within that masked patch. This procedure ensures that $Z_{gt}$ is inherently pixel-wise aligned with $Z_{f}$. The MAE is then the average of the absolute difference in depth values between the two maps. All reported MAE values are computed only on the sparse measurement points output by our method. Pixels without a valid depth estimate (shown as white in the visualized point clouds in Section VI) are excluded from the error computation.
@@ V-B Calibration Cross-Validation and Evaluation
A potential concern from the calibration performed, as described in Section IV-C, is that the optimal fusion weights determined using only the sphere geometry might be biased and not generalize well to other object shapes (outlined in detail in Section V). To investigate this, we performed a cross-validation analysis. For calibration, the sensor is rolled over the surface of the calibration sphere/object (rather than pressing the sphere into a static roller as is done typically with VBTS/NVBTS) while recording event data. From each rolling pass, a point cloud is generated and the optimization in Eq. 8 is performed within the masked contact region to determine the optimal parameters. We first determined the object-specific optimal fusion parameters $\theta_{i}$ (the weights $w_{s},w_{m},w_{e}$ were highlighted since the thresholding parameters did not change across objects) by running the grid search optimization (Eq. 8) using the ground truth $Z_{gt}^{(i)}$ for each object $i$ (where $i\in\{\text{Sphere, Stairs, Sparse Sphere}\}$).
@@
$e_{i}=\frac{1}{M}\sum_{i=1}^{M}\bigl\|Z_{f}^{(i)}(\theta_{i})-Z_{gt}^{(i)}\bigr\|_{1}$ (11)
@@ VI-A Single Object Reconstruction
As seen in Table III, the BMA fusion framework outperforms EMVS with an average percentage decrease in MAE of 25.2% across objects tested within the calibration set. Additionally, the fusion strategy of only fusing the left and right images that performed the best in reconstruction variability with Li et al. [li_2025_cyclic] was tested. While their method performs better than using the depth map generated using baseline EMVS, the proposed BMA fusion framework still outperforms the other methods. A quantitative analysis on a line array (three identical rectangles with height 0.6 mm) further confirms the improvement in depth consistency: the standard deviation of the reconstructed depth across the line array decreased significantly from 0.067 mm for the baseline EMVS reconstruction to 0.028 mm when using the BMA fusion approach, indicating a much more uniform depth estimation.
@@
BMA also generated depth maps with lower MAE across varying speeds as illustrated in Figure 9 in comparison to EMVS and Left & Right Fusion. However, the reduction in the MAE of the sphere is more significant at higher speeds, with a 5.27% reduction at 0.05 m/s, a 25.2% reduction at 0.3 m/s and starts to plateau with a 26.4% reduction at 0.5 m/s in comparison to EMVS. Interestingly, Figure 9 also shows that at lower speeds (e.g., 0.05 m/s), the performance difference between the BMA fusion (using all three reference views) and the Left & Right fusion (using only start and end views) becomes negligible. This suggests that while BMA fusion provides substantial benefits at higher speeds by effectively combining information from more distinct viewpoints to mitigate motion artifacts and improve consistency, at lower speeds the depth maps generated at the start, middle, and end times are likely very similar due to the minimal camera displacement within the time window. Consequently, fusing only the start and end views captures most of the available information, and adding the highly correlated midpoint view offers little additional advantage. These findings are in line with Bhattacharya et al. [bhattacharya2024monocular] as their event-based depth estimation model exhibits greater depth prediction quality at higher speeds. This effect could be explained by the fact that more events are generated within the same time window at higher speeds allowing for more accurate depth estimation through EMVS.
Another important consideration when it comes to BMA is the optimal number of reference views for fusion. To investigate this, we conducted an additional experiment on the sphere object using five reference views. This approach yielded a higher MAE of 0.1115 ± 0.0065 mm compared to the 0.0946 ± 0.0081 mm from the three-reference fusion, albeit with lower variability. This suggests a trade-off where adding more views can increase robustness but may also introduce additional warping errors that degrade overall accuracy. This finding is consistent with observations by Li et al. [li_2025_cyclic], who noted that excessive fusion can weaken unique information from individual viewpoints. The optimal number of reference views is likely task-dependent and differs depending on speed used, but for the conditions in our experiments, three views provided the best balance of accuracy and consistency.
@@
Object Average MAE (mm) Std. Dev. (mm) Sparse Sphere 0.0120 0.0010 Stairs 0.0581 0.0026 Bumps 0.0589 0.0024 Pyramid 0.0677 0.0058 Hexagon 0.0808 0.0042 Sphere 0.0946 0.0081 Overall 0.0620 0.0047
@@
To further assess the sensor’s qualitative performance and generalization, we tested it on three additional objects with diverse materials and surface complexities, as shown in Figure 10. The scanned objects include a rivet head with a surface defect, an M2 screw with fine threads, and a textured fabric. The resulting point clouds in each case exhibit strong geometric consistency with the physical objects. These results demonstrate the sensor’s robustness in reconstructing a variety of complex geometries beyond the tested 3D printed shapes and potential for industrial surface inspection applications.
Figure 10: Qualitative reconstruction results. (a) Physical objects: a rivet head with a defect, an M2 screw, and a textured fabric. (b) Corresponding 3D point cloud reconstructions. Color represents change in depth using JET colormap
@@ VI-B Large Surface Reconstruction
Figure 11: Examples of large surface reconstruction. (a) 3D printed varying-height line array (top) and its corresponding reconstruction (bottom). (b) 3D printed letter plate with varying-height letters and induced defects (top) and its reconstruction (bottom).
@@
The large surface reconstruction capabilities of our roller were tested with two 3D printed plates, shown in Figure 11. The final point cloud for each is generated by aligning results from multiple depth maps as done by Zhang et al. [zhang_gelroller_2024]. The multiple depth maps are aligned using ICP; since the speed and hence the translation of the roller are known from the robotic controller, the translation along the rolling direction is used as the initial pose for ICP. It is important to note that due to the high repeatability of the UR10 robot and the short travel distances in our experiments, the initial pose derived from proprioception is already highly accurate. While the ICP step provides minimal refinement in these specific tests, it is retained as a standard practice consistent with prior work [zhang_gelroller_2024] to ensure the pipeline’s robustness and generalizability for larger-area scans where robot drift could become a factor. The robot’s motion error within a single 20 ms reconstruction window is negligible compared to other error sources and does not significantly impact the precision of an individual depth map.
@@
The test objects’ reconstruction accuracy was averaged over five trials as was done for the single object reconstruction. The first test object was a line array with varying heights (0.1-1.2 mm at intervals of 0.1 mm), for which our sensor achieved a resulting MAE of 0.0638 $\pm$ 0.0055 mm across the whole surface (see Figure 11a). The second object was a letter plate, designed to benchmark against prior work in letter reconstruction like GelRoller [zhang_gelroller_2024] and to test defect detection (see Figure 11b). The letters on this plate have small, varying heights (ranging from 0.4 mm to 0.6 mm), and the plate includes both intentionally modeled and induced defects like scratches and printing artifacts. For the letter plate, our sensor achieved an MAE of 0.0596 ± 0.0035 mm, and as shown qualitatively in Figure 11b, it successfully captured the letter geometries. Also, the fine defects on the surface were also captured.
@@
As aforementioned, speed is critical for large surface reconstruction using tactile sensing. Our roller exhibits, to the best of our knowledge, the highest measured speed at 0.5 m/s or 500 mm/s for tactile scanning with an MAE below 100 microns. This is 11 times faster than the maximum reported speed for rolling belt tactile sensors (45 mm/s) [mirzaee_gelbelt_2025] and 45 times faster than the highest speed for roller tactile sensors (11 mm/s) [Cao2023].
@@
Table V summarizes the performance trade-offs for different continuous tactile sensing approaches based on the provided documents. Our roller significantly advances the state-of-the-art in terms of speed, achieving 500 mm/s while maintaining a competitive 3D reconstruction accuracy (62.0 $\mu m$ MAE). GelBelt demonstrates the potential for very high accuracy ($\sim$13.5 $\mu m$ average MAE, measured against a high-resolution GelSight Max scan as ground truth for specific defects) but operates at a considerably lower maximum speed of 45 mm/s [mirzaee_gelbelt_2025]. The GelRoller paper focused on a novel self-supervised reconstruction method, achieving 175 µm MAE in single-frame tests, but did not report a maximum scanning speed [zhang_gelroller_2024]. TouchRoller operated at the slowest speed (11 mm/s) and was designed for 2D texture mapping, not 3D reconstruction [Cao2023].
@@ VI-C High-Speed Tactile Braille Reading
The large surface tactile scanning of the roller was also tested with regards to feature detection and recognition. Two braille plates were scanned with standard braille sizing: semi sphere of radius of 0.635 mm and spacing of 2.54 mm. To obtain the Braille character recognition results, we processed the generated IWE, $I_{warp}$, (from Section IV-D) with an application-specific pipeline. From $I_{warp}$, we select a Region of Interest (ROI) corresponding to a single Braille cell location. This ROI, illustrated in Figure 12b, then undergoes an image processing pipeline to enhance dot visibility and reduce noise. The steps include: applying a bilateral filter for edge-preserving smoothing, using Contrast Limited Adaptive Histogram Equalization (CLAHE) to increase local contrast, performing binary thresholding to segment potential dots, and applying morphological opening and closing operations to eliminate salt-and-pepper noise.
@@
The processing pipeline above was utilized to process the events into a frame and subsequently detect the letter in the frame. Each frame is then stitched to the next frame to form the complete words (as seen in Figure 1 and Figure 13). Due to the high temporal resolution (microsecond resolution) of the event camera, the recognition algorithm performs well at low and high speeds (0.05 to 0.5 m/s). For comparison, Potdar et al. [potdar_high-speed_2024] demonstrated Braille reading at approximately 0.18 m/s with 87.5% accuracy. Our system achieves the same task at 0.5 m/s, which is 2.6 times faster, while maintaining full accuracy. Furthermore, as Potdar et al. highlight, this tactile feature recognition pipeline can be extended for dynamic detection of surface textures [potdar_high-speed_2024].
@@ VII Discussion And Limitations
Despite the promising performance, the current approach exhibits limitations. The depth maps generated by the EMVS algorithm, relying on ray-counting from asynchronous and sparse event streams, are inherently semi-dense. While this works well for reconstructing objects with distinct edges or sufficient texture to generate events, it can struggle with accurately capturing smooth, continuous surface geometries where event generation might be sparse, potentially leading to incomplete or interpolated reconstructions. Moreover, the event camera used, the DVXplorer mini, has a resolution of 640x480 pixels (VGA). This inherently limits the spatial resolution achievable by the roller tactile sensor. Currently, event camera technology is less mature than conventional camera technology. As the field advances, future event cameras with higher resolutions are expected. Integrating these improved cameras would significantly enhance the spatial resolution of our sensor design.

### v3

@@ V-A Experimental Setup
The subsequent Results section (Sec. VI) details the outcomes of these experiments. The evaluations focus on: (1) 3D reconstruction accuracy, primarily quantified using Mean Absolute Error (MAE) against ground truth geometries, including an analysis of the proposed BMA fusion strategy; (2) the relationship between rolling speed and reconstruction accuracy; (3) the effectiveness of large surface reconstruction through measurement stitching; and (4) the speed and accuracy achieved during high-speed Braille character recognition. To quantify reconstruction accuracy, we calculate the Mean Absolute Error (MAE) between the reconstructed depth map, $Z_{f}$, and the ground truth, $Z_{\text{gt}}$. Due to the stochastic and sparse nature of event generation [7], a fixed dense ground truth grid cannot be used. Instead, we perform a sparse pixel-wise evaluation: for every reconstructed point at coordinates $(x,y)$ falling within the manually segmented contact region mask, the ideal depth $Z_{\text{gt}}(x,y)$ is computed analytically using the known object geometry. This ensures that $Z_{\text{gt}}$ is inherently pixel-wise aligned with $Z_{f}$, isolating depth estimation error from planar motion artifacts. The MAE is then computed as the average absolute difference for all valid sparse points within the mask. Notably, noise within the contact region is included in the error metric, while off-contact background noise (outside the contact region mask is excluded. All reported MAE values are computed only on the sparse measurement points output by our method. Pixels without a valid depth estimate (shown as white in the visualized point clouds in Section VI) are excluded from the error computation.
@@ V-B Calibration Cross-Validation and Evaluation
A potential concern from the calibration performed, as described in Section IV-C, is that the optimal fusion weights determined using only the sphere geometry might be biased and not generalize well to other object shapes (outlined in detail in Section V). To investigate this, we performed a cross-validation analysis. For calibration, the sensor is rolled over the surface of the calibration sphere/object (rather than pressing the sphere into a static roller as is done typically with VBTS/NVBTS) while recording event data. From each rolling pass, a point cloud is generated and the optimization in Eq. 8 is performed within the masked contact region to determine the optimal parameters. We first determined the object-specific optimal fusion parameters $\theta_{i}$ (the weights $w_{s},w_{m},w_{e}$ were highlighted since the thresholding parameters did not change across objects) by running the grid search optimization (Eq. 8) using the ground truth $Z_{\text{gt}}^{(i)}$ for each object $i$ (where $i\in\{\text{Sphere, Stairs, Sparse Sphere}\}$).
@@
$e_{i}=\frac{1}{M}\sum_{i=1}^{M}\bigl\|Z_{f}^{(i)}(\theta_{i})-Z_{\text{gt}}^{(i)}\bigr\|_{1}$ (11)
@@ VI-A Single Object Reconstruction
As seen in Table III, the BMA fusion framework outperforms EMVS with an average percentage decrease in MAE of 25.2% across objects tested within the calibration set. Additionally, the fusion strategy of only fusing the left and right images that performed the best in reconstruction variability with Li et al. [31] was tested. While their method performs better than using the depth map generated using baseline EMVS, the proposed BMA fusion framework still outperforms the other methods. A quantitative analysis on a line array (three identical rectangles with height 0.6 mm) further confirms the improvement in depth consistency: the standard deviation of the reconstructed depth across the line array decreased significantly from 0.067 mm for the baseline EMVS reconstruction to 0.028 mm when using the BMA fusion approach, indicating a much more uniform depth estimation.
@@
BMA also generated depth maps with lower MAE across varying speeds as illustrated in Figure 9 in comparison to EMVS and Left & Right Fusion. However, the reduction in the MAE of the sphere is more significant at higher speeds, with a 5.27% reduction at 0.05 m/s, a 25.2% reduction at 0.3 m/s and starts to plateau with a 26.4% reduction at 0.5 m/s in comparison to EMVS. Interestingly, Figure 9 also shows that at lower speeds (e.g., 0.05 m/s), the performance difference between the BMA fusion (using all three reference views) and the Left & Right fusion (using only start and end views) becomes negligible. This suggests that while BMA fusion provides substantial benefits at higher speeds by effectively combining information from more distinct viewpoints to mitigate motion artifacts and improve consistency, at lower speeds the depth maps generated at the start, middle, and end times are likely very similar due to the minimal camera displacement within the time window. Consequently, fusing only the start and end views captures most of the available information, and adding the highly correlated midpoint view offers little additional advantage. These findings are in line with Bhattacharya et al. [8] as their event-based depth estimation model exhibits greater depth prediction quality at higher speeds. This effect could be explained by the fact that more events are generated within the same time window at higher speeds allowing for more accurate depth estimation through EMVS.
Another important consideration when it comes to BMA is the optimal number of reference views for fusion. To investigate this, we conducted an additional experiment on the sphere object using five reference views. This approach yielded a higher MAE of 0.1115 ± 0.0065 mm compared to the 0.0946 ± 0.0081 mm from the three-reference fusion, albeit with lower variability. This suggests a trade-off where adding more views can increase robustness but may also introduce additional warping errors that degrade overall accuracy. This finding is consistent with observations by Li et al. [31], who noted that excessive fusion can weaken unique information from individual viewpoints. The optimal number of reference views is likely task-dependent and differs depending on speed used, but for the conditions in our experiments, three views provided the best balance of accuracy and consistency.
@@
Object Average MAE (mm) Std. Dev. (mm)
Sparse Sphere 0.0120 0.0010
Stairs 0.0581 0.0026
Bumps 0.0589 0.0024
Pyramid 0.0677 0.0058
Hexagon 0.0808 0.0042
Sphere 0.0946 0.0081
Overall 0.0620 0.0047
@@
To further assess the sensor’s qualitative performance and generalization, we tested it on four additional objects with diverse materials and surface complexities, as shown in Figure 10. The scanned objects include a healthy flush rivet, a defective rivet with a bent head, an M2 screw with fine threads, and a textured fabric. Comparing the healthy (Fig. 10 - Column 1) and defective (Fig. 10 - Column 2) rivets clearly demonstrates the sensor’s defect detection capability: the healthy rivet yields a uniform depth map indicating flushness, while the defective rivet reveals a distinct height deviation corresponding to the bent head. The resulting point clouds in all cases exhibit strong geometric consistency with the physical objects. These results demonstrate the sensor’s robustness in reconstructing a variety of complex geometries beyond the tested 3D printed shapes and potential for industrial surface inspection applications.
Figure 10: Qualitative reconstruction results. (a) Physical objects: a flush rivet head, a rivet head with a defect, an M2 screw, and a textured fabric. (b) Corresponding 3D point cloud reconstructions. Color represents change in depth using JET colormap
@@ VI-B Large Surface Reconstruction
Figure 11: Examples of large surface reconstruction. (a) 3D printed varying-height line array (top) and its corresponding reconstruction (bottom). (b) 3D printed letter plate with varying-height letters and induced defects (top) and its reconstruction (bottom). Zoomed insets highlight manually induced defects either with scratches or intentionally modeled fault as well as examples of a false positive and printing artifacts.
@@
The large surface reconstruction capabilities of our roller were tested with two 3D printed plates, shown in Figure 11. The final point cloud for each is generated by aligning results from multiple depth maps as done by Zhang et al. [62]. The multiple depth maps are aligned using ICP; since the speed and hence the translation of the roller are known from the robotic controller, the translation along the rolling direction is used as the initial pose for ICP. It is important to note that due to the high repeatability of the UR10 robot and the short travel distances in our experiments, the initial pose derived from proprioception is already highly accurate. Quantitatively, the ICP step provided a maximum MAE reduction of approximately 0.0002 mm compared to using proprioception alone in our tests. While this refinement is minimal, the step is retained as a standard practice consistent with prior work [62] to ensure the pipeline’s robustness and generalizability for larger-area scans where robot drift could become a factor. The robot’s motion error within a single 20 ms reconstruction window is negligible compared to other error sources and does not significantly impact the precision of an individual depth map.
@@
The test objects’ reconstruction accuracy was averaged over five trials as was done for the single object reconstruction. The first test object was a line array with varying heights (0.1-1.2 mm at intervals of 0.1 mm), for which our sensor achieved a resulting MAE of 0.0638 $\pm$ 0.0055 mm across the whole surface (see Figure 11a). The second object was a letter plate, designed to benchmark against prior work in letter reconstruction like GelRoller [62] and to test defect detection (see Figure 11b). The letters on this plate have small, varying heights (ranging from 0.4 mm to 0.6 mm). Crucially, we focused our validation on specific induced defects: manually added scratches (seen on ’I’ and the first ’E’, and on top of the ’T’) and a modeled fault added to the CAD design (on the side of the ’T’). As shown in the zoomed insets of Figure 11b, our sensor successfully captured these induced defects. Additionally, the sensor detected unintentional printing artifacts, such as the diagonal layer lines on top of the second ’E’ and on top of the ’O’. Notably, these artifacts were only reconstructed where their physical depth exceeded the sensor’s sensitivity threshold; shallower printing lines in other regions did not deform the elastomer sufficiently to be detected, causing them to be naturally filtered out. Moreover, due to the noisy nature of event generation, some false positives are observed in the point cloud that could be misinterpreted as defects (e.g: right of second ’E’). Further discussion regarding false positive mitigation and detection of shallow tactile features is present in Section VII. For the letter plate, our sensor achieved an MAE of 0.0596 ± 0.0035 mm.
@@
As aforementioned, speed is critical for large surface reconstruction using tactile sensing. Our roller exhibits, to the best of our knowledge, the highest measured speed at 0.5 m/s or 500 mm/s for tactile scanning with an MAE below 100 microns. This is 11 times faster than the maximum reported speed for rolling belt tactile sensors (45 mm/s) [35] and 45 times faster than the highest speed for roller tactile sensors (11 mm/s) [10].
@@
Table V summarizes the performance trade-offs for different continuous tactile sensing approaches based on the provided documents. Our roller significantly advances the state-of-the-art in terms of speed, achieving 500 mm/s while maintaining a competitive 3D reconstruction accuracy (62.0 $\mu m$ MAE). GelBelt demonstrates the potential for very high accuracy ($\sim$13.5 $\mu m$ average MAE, measured against a high-resolution GelSight Max scan as ground truth for specific defects) but operates at a considerably lower maximum speed of 45 mm/s [35]. The GelRoller paper focused on a novel self-supervised reconstruction method, achieving 175 µm MAE in single-frame tests, but did not report a maximum scanning speed [62]. TouchRoller operated at the slowest speed (11 mm/s) and was designed for 2D texture mapping, not 3D reconstruction [10].
@@ VI-C High-Speed Tactile Braille Reading
The large surface tactile scanning of the roller was also tested with regards to feature detection and recognition. Two braille plates were scanned with standard braille sizing: semi sphere of radius of 0.635 mm and spacing of 2.54 mm. To obtain the Braille character recognition results, we processed the generated IWE, $I_{\text{warp}}$, (from Section IV-D) with an application-specific pipeline. From $I_{\text{warp}}$, we select a Region of Interest (ROI) corresponding to a single Braille cell location. This ROI, illustrated in Figure 12b, then undergoes an image processing pipeline to enhance dot visibility and reduce noise. The steps include: applying a bilateral filter for edge-preserving smoothing, using Contrast Limited Adaptive Histogram Equalization (CLAHE) to increase local contrast, performing binary thresholding to segment potential dots, and applying morphological opening and closing operations to eliminate salt-and-pepper noise.
@@
The processing pipeline above was utilized to process the events into a frame and subsequently detect the letter in the frame. Each frame is then stitched to the next frame to form the complete words (as seen in Figure 1 and Figure 13). Due to the high temporal resolution (microsecond resolution) of the event camera, the recognition algorithm performs well at low and high speeds (0.05 to 0.5 m/s). For comparison, Potdar et al. [38] demonstrated Braille reading at approximately 0.18 m/s with 87.5% accuracy. Similarly, Xu et al. [56] utilized Spiking Neural Networks (SNNs) to achieve 95.33% accuracy in reliable Braille reading, though at a lower speed of 0.1 m/s. Our system achieves the same task at 0.5 m/s, which is 2.6 times faster, while maintaining full accuracy. Furthermore, as Potdar et al. highlight, this tactile feature recognition pipeline can be extended for dynamic detection of surface textures [38].
@@ VII Discussion And Limitations
Despite the promising performance, the current approach exhibits limitations. The depth maps generated by the EMVS algorithm, relying on ray-counting from asynchronous and sparse event streams, are inherently semi-dense. While this works well for reconstructing objects with distinct edges or sufficient texture to generate events, it can struggle with accurately capturing smooth, continuous surface geometries where event generation might be sparse, potentially leading to incomplete or interpolated reconstructions. Additionally, edges aligned parallel to the rolling direction may generate fewer events due to minimal temporal contrast change [36], resulting in minor reconstruction gaps at those specific orientations (e.g., the top/bottom edges of the rivet in Fig. 10). This effect is less apparent in large surface reconstructions, where slight kinematic variations over longer rolling distances reduce the likelihood of edges remaining perfectly parallel to the motion vector. Moreover, the event camera used, the DVXplorer mini, has a resolution of 640x480 pixels (VGA). This inherently limits the spatial resolution that can be achieved by the tactile roller sensor. Currently, event camera technology is less mature than conventional camera technology. However, higher resolution sensors are emerging, such as the 1-megapixel event sensor presented by Finateu et al. [16]. Although these sensors are currently larger than the DVXplorer mini used here, integrating such improved hardware in future iterations would significantly enhance the spatial resolution of our sensor design.
Regarding the large surface reconstruction in Section VI-B, while the system successfully detected induced defects, distinguishing true defects from sensor noise (false positives) remains a challenge common to high-sensitivity tactile sensors. In Fig. 11b, certain ”defect-like” patterns are caused by signal noise. In future applications, these can be mitigated using spatial filtering based on the metric nature of our reconstruction; for instance, applying a size threshold to disregard unconnected depth clusters smaller than a specific physical dimension (e.g., <0.2<0.2 mm) relevant to the inspection tolerance. Furthermore, the detection threshold is physically coupled to the elastomer properties. In future applications, this sensitivity can be tuned by adjusting the skin hardness such as using softer skins to detect micro-features, or harder skins to filter out negligible surface roughness, allowing the sensor response to be matched to the desired defect definition.